A->B

The rate of the reaction is 1.6 x10^-2 M/s when the concentration of A is 0.35M. Calculate the rate constant if the reaction is

a) first order in A
b) second order in A
1
Expert's answer
2020-06-29T07:44:04-0400

First order reaction:

Rate = k(A),


Where (k) is a constant


R = [ ( 1.6)(10-2)](0.35)


R = (5.6)(10-2)ms-1


R = 0.056ms-1


Second order reaction:


Rate = k ( A) 2


R = [(1.6)(10)-2 ](0.35)2


R = (1.96)(10-3)m-1s-1


R = 0.00196m-1s-1-